OpenStudy (anonymous):

Renee needs to burn at least 300 calories at the gym today. She can burn 25 calories per minute on the treadmill (t) and 10 calories per minute on the elliptical trainer (e). Which inequality below can be used to find the possible combinations of time she could spend on each machine and still meet her goal of burning at least 300 calories? 25t + 10e 300 25t - 10e 300 25t + 10e 300 25t - 10e 300

OpenStudy (anonymous):

25t + 10e >= 300 25t - 10e >= 300 25t + 10e <= 300 25t - 10e <= 300

OpenStudy (anonymous):

@dmezzullo

OpenStudy (anonymous):

@churrity

OpenStudy (anonymous):

@poopsiedoodle

OpenStudy (poopsiedoodle):

It's A, because you're not going to subtract anything, and you need 300 or more calories, not less than 300.

OpenStudy (anonymous):

thank you!!

OpenStudy (poopsiedoodle):